# Constrained free energy minimization for the design of thermal states and stabilizer thermodynamic systems

### v1

@@ I.1 Background and motivation
Beyond establishing a link between quantum thermodynamics and semi-definite optimization, Ref. [16] proposed various classical and hybrid quantum-classical (HQC) algorithms for solving semi-definite programs, being motivated by physical intuition coming from quantum thermodynamics when applied to the special class of constrained energy minimization problems. These algorithms, hereafter referred to as the LMPW25 algorithms, have provable guarantees for their runtimes and are based on the observation that parameterized thermal states, also known as non-Abelian thermal states [22, 23, 24] or quantum Boltzmann machines [25, 26, 27], are optimal for solving constrained free-energy minimization problems, which closely approximate the original constrained energy minimization problems for sufficiently low temperatures. Indeed, Ref. [16] established that the optimization landscapes of dual chemical potential maximization problems are concave in the parameters of the thermal states, so that conventional optimization techniques like gradient ascent, stochastic gradient ascent, and their variants [28] are guaranteed to converge. The HQC algorithms are variational in nature, using a quantum computer for two purposes only, which include the preparation of parameterized thermal states and estimating expectations of observables (i.e., the Hamiltonian and non-commuting charges). A classical computer is then employed for storing and updating the parameters of the parameterized thermal states as the HQC algorithm progresses toward convergence. The LMPW25 HQC algorithms make use of first-derivative information to determine the next step in a search and can additionally incorporate second-derivative information to improve their convergence to globally optimal solutions.
@@
Let us also recall an observation from [16], that the LMPW25 HQC algorithms for solving constrained free energy minimization problems can alternatively be viewed as HQC algorithms for preparing non-Abelian thermal states of thermodynamic systems with conserved, non-commuting charges. This aspect of the LMPW25 HQC algorithms could be useful in fields like materials science, condensed matter physics, and quantum chemistry, because one expects such states to arise in these thermodynamic systems and, after preparing them, one could subsequently measure various observables of interest on the generated states, in order to determine their properties.
@@ I.2 Summary of contributions
In this paper, we apply the LMPW25 algorithms to concrete examples of constrained energy minimization problems of interest in quantum thermodynamics, including the one- and two-dimensional Heisenberg models [37] with nearest and next-to-nearest neighbor interactions, and with the non-commuting charges set to the total magnetizations in the $x$, $y$, and $z$ directions. Such models describe interacting spin systems on a lattice and are essential for understanding magnetic materials [38].
As a first conceptual contribution, we offer an alternative compelling interpretation of the LMPW25 algorithms as methods for designing ground and thermal states of controllable Hamiltonians, such that these states satisfy desired properties (i.e., constraints on the expectation values of certain observables). This application of the algorithms could find use in the design of molecules and materials with desired properties.
@@
Moreover, we utilize this correspondence to demonstrate that the LMPW25 HQC algorithms can serve as thermodynamics-inspired HQC algorithms for encoding quantum information into quantum error-correcting codes, which is rather different from the standard approach that employs an encoding circuit to do so (see, e.g., [56, Chapter 4] and [57, Section 3.5]). The approach is instead based on constrained energy minimization and the fact that the LMPW25 HQC algorithms, as applied in this context, prepare low-temperature thermal states of stabilizer thermodynamic systems. We suspect that this approach will be useful in experimental scenarios for which it is accessible to prepare low-temperature thermal states of stabilizer thermodynamic systems.
As further examples of constrained energy minimization, and building on the aforementioned bridge, we also benchmark the LMPW25 algorithms on stabilizer thermodynamic systems constructed from the one-to-three-qubit repetition code [58], the perfect one-to-five-qubit code [59, 60], and the two-to-four qubit error-detecting code [61, 62]. For each case, we demonstrate that the LMPW25 HQC algorithms are capable of steering the state of the stabilizer thermodynamic system in such a way that any desired quantum state is encoded into the code’s logical qubits approximately, with the approximation improving as the temperature of the system decreases.
On the one hand, the numerical contributions of our paper can be viewed as complementing the theoretical developments of [16], by providing a comprehensive numerical study of the performance of the LMPW25 algorithms for constrained energy minimization problems. The value of this contribution is that it offers a practical assessment of the theoretical guarantees from [16] for various problems of interest in quantum thermodynamics. On the other hand, the first conceptual contribution of our paper, i.e., the design of ground and thermal states, could have far reaching applications for the use of quantum computers in molecular and material design. Additionally, the other theoretical contribution of our paper, i.e., establishing a bridge between quantum error correction and conserved, non-commuting charges in quantum thermodynamics via the notion of stabilizer thermodynamic systems, is conceptually enlightening and the related encoding algorithm could be useful as an alternative approach for stabilizing quantum information.
@@ I.3 Overview of numerical simulations
Before proceeding with the details of our paper, let us first elaborate on some aspects of the numerical simulations that we conducted. Given that all of the constrained energy minimization problems we consider can be formulated as semi-definite programs (SDPs) [63, Section 4.6.2], the performance of the LMPW25 algorithms can be benchmarked against the values returned by standard SDP solvers. These solvers are based on the interior-point method [64, 65, 66] and are so reliable that we can view their outputs as essentially the true optimal values of the constrained energy minimization problems. We test out four variants of the LMPW25 algorithms, which include a first-order classical algorithm, a second-order classical algorithm, a first-order HQC algorithm, and a second-order HQC algorithm. Considering the classical algorithms in addition to the HQC algorithms is useful, as the classical algorithms can be understood as fully noiseless simulators of the HQC algorithms, such that the only kind of error involved when using them is the algorithmic error and not any other kind of error, e.g., shot noise arising from estimating expectations of observables. For this reason, we can benchmark the HQC algorithms against the classical algorithms, in addition to the true values generated by standard SDP solvers.
We conducted all of our simulations for small system sizes, between three to six qubits. These sizes are reasonable for providing a sense of the performance of the LMPW25 algorithms, while taking a reasonable amount of time for them to converge when conducted on readily available laptop computers. In order to speed up the second-order HQC algorithms, which require estimating the Hessian, we parallelized the implementation of them in order to use multiple processor cores. All of the software code needed to reproduce the results of our paper is publicly available as a zenodo repository https://zenodo.org/records/16814266. We leave it to future research to scale up the examples to much larger numbers of qubits, possibly by using tensor-network methods [67, 68, 69], and to conduct the simulations on actual quantum computers, incorporating techniques like error correction [41] and mitigation [70] in order to reduce the effects of noise.
@@ I.3.1 Method for thermal state preparation
As mentioned in Section I.1, a critical subroutine used by the first- and second-order HQC algorithms is thermal state preparation, and there has been much development on this topic in recent years [29, 30, 31, 32, 33, 34, 35, 36]. However, in our paper, we do not make use of these more recent advanced methods. Instead, we employ a simple method for thermal state preparation, consisting of feeding the numerical values of the desired thermal state to Qiskit’s DensityMatrix routine [71] and employing its functionality for generic state preparation when given a description of the state. Although this method is elementary, it has value in demonstrating the performance of the algorithms under the assumption that a robust method for thermal state preparation exists. Indeed, the literature on thermal-state preparation algorithms is rather active and advances are being made rapidly, such that we expect continued progress in this direction during the near term. As such, any of these alternative thermal state preparation algorithms can be plugged in to our approach, and further study of the LMPW25 HQC algorithms’ performance can be conducted under them.
@@ I.4 Paper organization
Our paper is organized as follows. Section II provides review material required for understanding the rest of the paper. In particular, Section II.1 reviews quantum thermodynamics in the presence of conserved, non-commuting charges, Section II.2 reviews the LMPW25 classical and HQC algorithms for constrained energy minimization, Section II.3 reviews quantum Heisenberg models, and Section II.4 reviews quantum error correction codes with a focus on stabilizer codes.
Our contributions start with Section III, where we offer an alternative interpretation of the LMPW25 algorithms as methods for designing Hamiltonians with ground and thermal states satisfying desired properties. In Section IV, we observe that every stabilizer code corresponds to a stabilizer thermodynamic system, in which the Hamiltonian is formed from the code’s stabilizer operators and the conserved, non-commuting charges are formed from the code’s logical operators. In Section V, we argue how the LMPW25 HQC algorithms for constrained energy minimization can be used to encode quantum information into stabilizer codes when applied to stabilizer thermodynamic systems. We then detail our simulation results in Section VI, exhibiting the performance of the LMPW25 classical and HQC algorithms on various quantum Heisenberg models and stabilizer thermodynamic systems. We finally conclude in Section VII with a summary of our findings and suggestions for future research.
@@ II.1 Quantum thermodynamics with conserved, non-commuting charges
All of the thermodynamic models that we deal with in this paper, including quantum Heisenberg models and stabilizer thermodynamic systems, satisfy the charge conservation constraint in (4). However, we should note that the developments in [16] do not depend on this constraint holding, nor does the interpretation of the LMPW25 algorithms given in Section III.
II.2 LMPW25 algorithms for constrained energy minimization
In this section, we review the problem of constrained energy minimization, as well as the LMPW25 classical and HQC algorithms for solving it [16].
@@ II.2.1 Review of constrained energy minimization
so that one has $\varepsilon\geq 0$ additive error in the approximation by picking $T=\frac{\varepsilon}{n\ln d}$. Observe that this choice implies that the temperature is low, i.e., proportional to the desired error $\varepsilon$ and inversely proportional to the number of qudits in the system.
@@
II.2.2 LMPW25 first-order classical algorithm
With the background from Section II.2.1 in hand, let us now recall the LMPW25 first-order classical algorithm for constrained energy minimization [16, Algorithm 1]:
@@ II.2.2 LMPW25 first-order classical algorithm
    Set $m\leftarrow 0$, and initialize $\mu^{m}\leftarrow\left(0,\ldots,0\right)$, the temperature $T\leftarrow\frac{\varepsilon}{n\ln d}$, the step size $\eta\in[0,L^{-1})$, where $L$ is defined in (17), and the stopping parameter $\delta>0$.
@@
so that the output of Algorithm 1 in (20) is an approximation of $f(\mu^{m})$ that omits the last term $TS(\rho_{T}(\mu))$ in (21). As stated around [16, Eq. (21)], we could alternatively output $f(\mu^{m})$ as an approximation of the desired solution $E(\mathcal{Q},q)$, but the output in (20) is consistent with the output of the LMPW25 HQC algorithms. For our purposes here, this allows for a more direct comparison of the classical first-order algorithm above and the first-order HQC algorithm recalled in Section II.2.3.
We have phrased the LMPW25 classical algorithm slightly differently from how it was presented previously in [16, Algorithm 1], in order to simplify its presentation. The main difference is the introduction of the stopping parameter $\delta$ above, as a simple way of ending the algorithm after convergence has been achieved (i.e., when the norm of the gradient is small, further changes to $\mu$ are small and convergence near a globally optimal solution is guaranteed at this point by [76, Corollary 3.5]).
@@
Before moving on to the LMPW25 first-order HQC algorithm reviewed in Section II.2.3, let us note that the LMPW25 first-order classical algorithm can be considered a noiseless simulation of the first-order HQC algorithm. The main reason for this is that the classical algorithm numerically computes the expectation values $\operatorname{Tr}[H\rho_{T}(\mu)]$ and $\operatorname{Tr}[Q_{i}\rho_{T}(\mu)]$, for all $i\in\left[c\right]$, whereas the HQC algorithm estimates these expectation values through sampling on a quantum computer. As such, the classical algorithm can serve as a benchmark for the HQC algorithm, in addition to the benchmark provided by standard SDP solvers.
II.2.3 LMPW25 first-order hybrid quantum-classical algorithm
@@
Algorithm 1 is then modified as follows, to become the LMPW25 first-order HQC algorithm for constrained energy minimization:
@@
    Set $m\leftarrow 0$, and initialize $\mu^{m}\leftarrow\left(0,\ldots,0\right)$, the temperature $T\leftarrow\frac{\varepsilon}{n\ln d}$, the step size $\eta\in[0,L^{-1})$, where $L$ is defined in (17), and the stopping parameter $\delta>0$.
@@
As presented above, Algorithm 2 is a simplification of the first-order HQC algorithm proposed in [16, Algorithm 2]. Since the gradient calculations are replaced by gradient estimates, the latter are random variables, and the algorithm thus employs stochastic gradient ascent in its search for a globally optimal solution (see, e.g., [76, Chapter 5] and [28, Section 6]). Since one of the goals of [16] was to prove that the algorithm converges to a global optimum, the presentation of the algorithm was more involved in order to provide a theoretical guarantee on its performance. However, in practice, we find in our simulations that the above simplification of the algorithm works well for converging near the global optimum.
@@
II.2.4 LMPW25 second-order classical algorithm

### v2

@@ I.1 Background and motivation
Beyond establishing a link between quantum thermodynamics and semi-definite optimization, Ref. [16] proposed various classical and hybrid quantum-classical (HQC) algorithms for solving semi-definite programs, being motivated by physical intuition coming from quantum thermodynamics when applied to the special class of constrained energy minimization problems. These algorithms, hereafter referred to as the LMPW algorithms, have provable guarantees for their runtimes and are based on the observation that parameterized thermal states, also known as non-Abelian thermal states [22, 23, 24] or quantum Boltzmann machines [25, 26, 27], are optimal for solving constrained free-energy minimization problems, which closely approximate the original constrained energy minimization problems for sufficiently low temperatures. Indeed, Ref. [16] established that the optimization landscapes of dual chemical potential maximization problems are concave in the parameters of the thermal states, so that conventional optimization techniques like gradient ascent, stochastic gradient ascent, and their variants [28] are guaranteed to converge. The HQC algorithms are variational in nature, using a quantum computer for two purposes only, which include the preparation of parameterized thermal states and estimating expectations of observables (i.e., the Hamiltonian and non-commuting charges). A classical computer is then employed for storing and updating the parameters of the parameterized thermal states as the HQC algorithm progresses toward convergence. The LMPW HQC algorithms make use of first-derivative information to determine the next step in a search and can additionally incorporate second-derivative information to improve their convergence to globally optimal solutions.
@@
Let us also recall an observation from [16], that the LMPW HQC algorithms for solving constrained free energy minimization problems can alternatively be viewed as HQC algorithms for preparing non-Abelian thermal states of thermodynamic systems with conserved, non-commuting charges. This aspect of the LMPW HQC algorithms could be useful in fields like materials science, condensed matter physics, and quantum chemistry, because one expects such states to arise in these thermodynamic systems and, after preparing them, one could subsequently measure various observables of interest on the generated states, in order to determine their properties.
@@ I.2 Summary of contributions
In this paper, we apply the LMPW algorithms to concrete examples of constrained energy minimization problems of interest in quantum thermodynamics, including the one- and two-dimensional Heisenberg models [37] with nearest and next-to-nearest neighbor interactions, and with the non-commuting charges set to the total magnetizations in the $x$, $y$, and $z$ directions. Such models describe interacting spin systems on a lattice and are essential for understanding magnetic materials [38].
As a first conceptual contribution, we offer an alternative compelling interpretation of the LMPW algorithms as methods for designing ground and thermal states of controllable Hamiltonians, such that these states satisfy desired properties (i.e., constraints on the expectation values of certain observables). This application of the algorithms could find use in the design of molecules and materials with desired properties.
@@
Moreover, we utilize this correspondence to demonstrate that the LMPW HQC algorithms can serve as thermodynamics-inspired HQC algorithms for encoding quantum information into quantum error-correcting codes, which is rather different from the standard approach that employs an encoding circuit to do so (see, e.g., [56, Chapter 4] and [57, Section 3.5]). The approach is instead based on constrained energy minimization and the fact that the LMPW HQC algorithms, as applied in this context, prepare low-temperature thermal states of stabilizer thermodynamic systems. We suspect that this approach will be useful in experimental scenarios for which it is accessible to prepare low-temperature thermal states of stabilizer thermodynamic systems.
As further examples of constrained energy minimization, and building on the aforementioned bridge, we also benchmark the LMPW algorithms on stabilizer thermodynamic systems constructed from the one-to-three-qubit repetition code [58], the perfect one-to-five-qubit code [59, 60], and the two-to-four qubit error-detecting code [61, 62]. For each case, we demonstrate that the LMPW HQC algorithms are capable of steering the state of the stabilizer thermodynamic system in such a way that any desired quantum state is encoded into the code’s logical qubits approximately, with the approximation improving as the temperature of the system decreases.
On the one hand, the numerical contributions of our paper can be viewed as complementing the theoretical developments of [16], by providing a comprehensive numerical study of the performance of the LMPW algorithms for constrained energy minimization problems. The value of this contribution is that it offers a practical assessment of the theoretical guarantees from [16] for various problems of interest in quantum thermodynamics. On the other hand, the first conceptual contribution of our paper, i.e., the design of ground and thermal states, could have far reaching applications for the use of quantum computers in molecular and material design. Additionally, the other theoretical contribution of our paper, i.e., establishing a bridge between quantum error correction and conserved, non-commuting charges in quantum thermodynamics via the notion of stabilizer thermodynamic systems, is conceptually enlightening and the related encoding algorithm could be useful as an alternative approach for stabilizing quantum information.
@@ I.3 Overview of numerical simulations
Before proceeding with the details of our paper, let us first elaborate on some aspects of the numerical simulations that we conducted. Given that all of the constrained energy minimization problems we consider can be formulated as semi-definite programs (SDPs) [63, Section 4.6.2], the performance of the LMPW algorithms can be benchmarked against the values returned by standard SDP solvers. These solvers are based on the interior-point method [64, 65, 66] and are so reliable that we can view their outputs as essentially the true optimal values of the constrained energy minimization problems. We test out four variants of the LMPW algorithms, which include a first-order classical algorithm, a second-order classical algorithm, a first-order HQC algorithm, and a second-order HQC algorithm. Considering the classical algorithms in addition to the HQC algorithms is useful, as the classical algorithms can be understood as fully noiseless simulators of the HQC algorithms, such that the only kind of error involved when using them is the algorithmic error and not any other kind of error, e.g., shot noise arising from estimating expectations of observables. For this reason, we can benchmark the HQC algorithms against the classical algorithms, in addition to the true values generated by standard SDP solvers.
We conducted all of our simulations for small system sizes, between three to six qubits. These sizes are reasonable for providing a sense of the performance of the LMPW algorithms, while taking a reasonable amount of time for them to converge when conducted on readily available laptop computers. In order to speed up the second-order HQC algorithms, which require estimating the Hessian, we parallelized the implementation of them in order to use multiple processor cores. All of the software code needed to reproduce the results of our paper is publicly available as a zenodo repository https://zenodo.org/records/17285628. We leave it to future research to scale up the examples to much larger numbers of qubits, possibly by using tensor-network methods [67, 68, 69], and to conduct the simulations on actual quantum computers, incorporating techniques like error correction [41] and mitigation [70] in order to reduce the effects of noise.
@@ I.3.1 Method for thermal state preparation
As mentioned in Section I.1, a critical subroutine used by the first- and second-order HQC algorithms is thermal state preparation, and there has been much development on this topic in recent years [29, 30, 31, 32, 33, 34, 35, 36]. However, in our paper, we do not make use of these more recent advanced methods. Instead, we employ a simple method for thermal state preparation, consisting of feeding the numerical values of the desired thermal state to Qiskit’s DensityMatrix routine [71] and employing its functionality for generic state preparation when given a description of the state. Although this method is elementary, it has value in demonstrating the performance of the algorithms under the assumption that a robust method for thermal state preparation exists. Indeed, the literature on thermal-state preparation algorithms is rather active and advances are being made rapidly, such that we expect continued progress in this direction during the near term. As such, any of these alternative thermal state preparation algorithms can be plugged in to our approach, and further study of the LMPW HQC algorithms’ performance can be conducted under them.
@@ I.4 Paper organization
Our paper is organized as follows. Section II provides review material required for understanding the rest of the paper. In particular, Section II.1 reviews quantum thermodynamics in the presence of conserved, non-commuting charges, Section II.2 reviews the LMPW classical and HQC algorithms for constrained energy minimization, Section II.3 reviews quantum Heisenberg models, and Section II.4 reviews quantum error correction codes with a focus on stabilizer codes.
Our contributions start with Section III, where we offer an alternative interpretation of the LMPW algorithms as methods for designing Hamiltonians with ground and thermal states satisfying desired properties. In Section IV, we observe that every stabilizer code corresponds to a stabilizer thermodynamic system, in which the Hamiltonian is formed from the code’s stabilizer operators and the conserved, non-commuting charges are formed from the code’s logical operators. In Section V, we argue how the LMPW HQC algorithms for constrained energy minimization can be used to encode quantum information into stabilizer codes when applied to stabilizer thermodynamic systems. We then detail our simulation results in Section VI, exhibiting the performance of the LMPW classical and HQC algorithms on various quantum Heisenberg models and stabilizer thermodynamic systems. We finally conclude in Section VII with a summary of our findings and suggestions for future research.
@@ II.1 Quantum thermodynamics with conserved, non-commuting charges
All of the thermodynamic models that we deal with in this paper, including quantum Heisenberg models and stabilizer thermodynamic systems, satisfy the charge conservation constraint in (4). However, we should note that the developments in [16] do not depend on this constraint holding, nor does the interpretation of the LMPW algorithms given in Section III.
II.2 LMPW algorithms for constrained energy minimization
In this section, we review the problem of constrained energy minimization, as well as the LMPW classical and HQC algorithms for solving it [16].
@@ II.2.1 Review of constrained energy minimization
so that one has $\varepsilon\geq 0$ additive error in the approximation by picking $T=\frac{\varepsilon}{n\ln d}$  [^1] Let us note that the dimension of the system we are considering here is $d^{n}$, instead of $d$ as considered previously in [16]. Observe that this choice implies that the temperature is low, i.e., proportional to the desired error $\varepsilon$ and inversely proportional to the number of qudits in the system.
@@
II.2.2 LMPW first-order classical algorithm
With the background from Section II.2.1 in hand, let us now recall the LMPW first-order classical algorithm for constrained energy minimization [16, Algorithm 1]:
@@ II.2.2 LMPW first-order classical algorithm
    Set $m\leftarrow 0$, and initialize $\mu^{m}\leftarrow\left(0,\ldots,0\right)$, the temperature $T\leftarrow\frac{\varepsilon}{n\ln d}$, the step size $\eta\in(0,L^{-1}]$, where $L$ is defined in (17), and the stopping parameter $\delta>0$.
@@
so that the output of Algorithm 1 in (20) is an approximation of $f(\mu^{m})$ that omits the last term $TS(\rho_{T}(\mu))$ in (21). As stated around [16, Eq. (21)], we could alternatively output $f(\mu^{m})$ as an approximation of the desired solution $E(\mathcal{Q},q)$, but the output in (20) is consistent with the output of the LMPW HQC algorithms. For our purposes here, this allows for a more direct comparison of the classical first-order algorithm above and the first-order HQC algorithm recalled in Section II.2.3.
We have phrased the LMPW classical algorithm slightly differently from how it was presented previously in [16, Algorithm 1], in order to simplify its presentation. The main difference is the introduction of the stopping parameter $\delta$ above, as a simple way of ending the algorithm after convergence has been achieved (i.e., when the norm of the gradient is small, further changes to $\mu$ are small and convergence near a globally optimal solution is guaranteed at this point by [77, Corollary 3.5]).
@@
Before moving on to the LMPW first-order HQC algorithm reviewed in Section II.2.3, let us note that the LMPW first-order classical algorithm can be considered a noiseless simulation of the first-order HQC algorithm. The main reason for this is that the classical algorithm numerically computes the expectation values $\operatorname{Tr}[H\rho_{T}(\mu)]$ and $\operatorname{Tr}[Q_{i}\rho_{T}(\mu)]$, for all $i\in\left[c\right]$, whereas the HQC algorithm estimates these expectation values through sampling on a quantum computer. As such, the classical algorithm can serve as a benchmark for the HQC algorithm, in addition to the benchmark provided by standard SDP solvers.
II.2.3 LMPW first-order hybrid quantum-classical algorithm
@@
Algorithm 1 is then modified as follows, to become the LMPW first-order HQC algorithm for constrained energy minimization:
@@
    Set $m\leftarrow 0$, and initialize $\mu^{m}\leftarrow\left(0,\ldots,0\right)$, the temperature $T\leftarrow\frac{\varepsilon}{n\ln d}$, the step size $\eta\in(0,L^{-1}]$, where $L$ is defined in (17), and the stopping parameter $\delta>0$.
@@
As presented above, Algorithm 2 is a simplification of the first-order HQC algorithm proposed in [16, Algorithm 2]. Since the gradient calculations are replaced by gradient estimates, the latter are random variables, and the algorithm thus employs stochastic gradient ascent in its search for a globally optimal solution (see, e.g., [77, Chapter 5] and [28, Section 6]). Since one of the goals of [16] was to prove that the algorithm converges to a global optimum, the presentation of the algorithm was more involved in order to provide a theoretical guarantee on its performance. However, in practice, we find in our simulations that the above simplification of the algorithm works well for converging near the global optimum.
@@
II.2.4 LMPW second-order classical algorithm